OpenStudy (anonymous):

Solve the following equations for x between 0 and 360 degrees (a) sinx + 4cos^2x = 1 (b) 2sec^2x = 3 - tan^2x For part (a) I have gotten this far: (4sinx + 3) (sinx - 1) = 0 How to I solve for x from there? Same applies to part (b): (tanx - 1) (tanx + 3) = 0. How to solve for x? Please show all working thanks.

OpenStudy (anonymous):

you said, sin(x) + 4(sin^2(x) - 1) = 1 isn't it supposed to be sin^2(x) + cos^2(x) = 1 so if you're gonna replace cos^2(x), it will be 1- Sin^2(x) not sin^2(x) - 1 that will be -cos^2(x) ?

hero (hero):

sin(x) + 4(1 - sin^2(x)) = 1 sin(x) + 4 - 4sin^2(x) = 1 sin(x) +3 - 4sin^2(x) = 0 4 sin^2(x) - sin(x) - 3 = 0 4 sin(x)^2 - sin(x) - 3 = 0 4x^2 - x - 3 = 0 4x^2 - 4x + 3x - 3 = 0 4x(x - 1) + 3(x - 1) = 0 (x - 1)(4x + 3) = 0 x - 1 = 0 4x + 3 = 0 sin(x) - 1 = 0 4(sin(x)) + 3 = 0 sin(x) = 1 sin(x) = -3/4 x = sin^{-1}(1) x = sin^{-1}(-3/4) x = pi/2 x = no solution

hero (hero):

@ambitiouschick